OpenStudy (anonymous):

The sum of the first 100 numbers, 1 to 100 is divisible by: a) 2,4,& 8 b) 2&4 c)2 only d)none of these

OpenStudy (anonymous):

c)2 only

OpenStudy (anonymous):

sum of numbers from 1 to 100 = (100/2)(1+100)=50*101=5050 i think 2 only will be ans ...as it is divisible by only 2

OpenStudy (anonymous):

good mr. prashant